# From Complex-Analytic Models to Sparse Domination: A Dyadic Approach of Hypersingular Operators via Bourgain’s Interpolation Method

### v4

@@ 2. Notations
The Lorentz space $L^{p,1}({\mathbb{D}})$ is defined by
$\|f\|_{L^{p,1}({\mathbb{D}})}:=p\int_{0}^{\infty}\lambda^{p-1}\big|\{z\in{\mathbb{D}}:\ |f(z)|>\lambda\}\big|\,d\lambda<\infty,$
@@ 3.1. Off-critical line estimate
For any $\varepsilon>0$, ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{\infty}({\mathbb{D}})\to L^{\frac{1}{2t-2+\varepsilon}}({\mathbb{D}})$ is bounded.
@@
Observe that it can happen that the intersection of the two sets of Carleson boxes $\{Q_{4^{tk},i}\}_{i\geq 1}$ and $\{Q_{4^{t(k+1)},i}\}_{i\geq 1}$ is not empty. Hence, for each $k\in\mathbb{Z}$, define the disjoint union of Carleson boxes
@@ 1. $\bullet$
    $E_{4^{tk}}\backslash E_{4^{t(k+1)}}\subseteq\bigcup_{i\geq 1}\widetilde{Q}_{4^{tk},i}$.
@@ 2. $\bullet$
    $\{\widetilde{Q}^{\textrm{up}}_{4^{tk},i}\}_{k\geq\mathbb{Z},\;i\geq 1}$ are mutually disjoint.
Using now (3.1) with $E_{4^{tk}}$ replaced by $E_{4^{tk}}\backslash E_{4^{t(k+1)}}$, and $\{Q_{4^{tk},i}\}_{i\geq 1}$ replaced by $\{\widetilde{Q}_{4^{tk},i}\}_{i\geq 1}$, respectively, we have
RHS of (3.3) $\leq\sum_{k\in\mathbb{Z}}\left(\sum_{i=1}^{\infty}\int_{\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)\right)^{3-2t-\varepsilon}\cdot\left(\sum_{i=1}^{\infty}|\widetilde{Q}_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}\right)^{2t+\varepsilon-2}$
(3.5) $\leq\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)\right)^{3-2t-\varepsilon}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|\widetilde{Q}_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}\right)^{2t+\varepsilon-2}$
Since $\{\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}\}_{k\in\mathbb{Z},\;i\geq 1}$ are mutually disjoint, therefore, the first double sum in (3.1) is bounded above by
@@
$\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|\widetilde{Q}_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}$ $\simeq\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}=\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\left|\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}\right|\cdot\frac{1}{\left|\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}\right|^{\frac{t-1}{2t+\varepsilon-2}}}$
$\simeq\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}}\frac{1}{(1-|z|^{2})^{\frac{2(t-1)}{2t+\varepsilon-2}}}dA(z)\lesssim\int_{{\mathbb{D}}}\frac{1}{(1-|z|^{2})^{\frac{2(t-1)}{2t+\varepsilon-2}}}dA(z)$
$=\int_{0}^{1}\frac{1}{(1-r)^{\frac{2(t-1)}{2t+\varepsilon-2}}}dr<+\infty.$
@@ 3.2. Critical line estimate
We first deal with the endpoint $\left(\frac{1}{p},\frac{1}{q}\right)=(0,2t-2)$.
@@
Sufficiency. Let $\alpha>0$ and denote $E:=\left\{z\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)>\alpha\right\}$. As usual, we decompose $E$ into a union of maximal and mutually disjoint Carleson boxes $\{Q_{i}\}_{i\geq 1}$ satisfying $Q_{i}=Q_{I_{i}}$ for $I_{i}\subset\mathbb{T}$ and
@@
then for every $0<r\leq\infty$, the operator ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ extends a bounded operator from $L^{p,r}({\mathbb{D}})$ to $L^{q,r}({\mathbb{D}})$. In particular, ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{p}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$ is bounded.
@@
The first assertion follows from the off-diagonal Marcinkiewicz interpolation theorem [16, Theorem 1.4.19]. The second follows from by letting $r=\infty$ and the fact that $L^{p}({\mathbb{D}})\subset L^{p,\infty}({\mathbb{D}})$. ∎
@@ 2. $\bullet$
Sufficiency. Again, for any $\alpha>0$, we denote $E_{\alpha}:=\left\{z\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)>\alpha\right\}$. Then as usual, one can decompose $E_{\alpha}$ into disjoint union of maximal Carleson boxes $\{Q_{\alpha,i}\}_{i\geq 1}$ enjoying (3.1). Then following the argument in (3.1) with applying Hölder for the conjugate pair $(1/l,1/l^{\prime})=(3-2t,2t-2)$, we have
@@
Then using (3.2) with again $E_{4^{tk}}$ replaced by $E_{4^{tk}}\backslash E_{4^{t(k+1)}}$, and $\{Q_{4^{tk},i}\}_{i\geq 1}$ replaced by $Q_{4^{tk},i}\}_{i\geq 1}\backslash\{Q_{4^{t(k+1)},i}\}_{i\geq 1}$, respectively, we have
@@
$\leq\sum_{k\in\mathbb{Z}}\left(\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textnormal{up}}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t}}\omega(z)dA(z)\right)^{3-2t}\left(\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}$
$\lesssim\left\|{\mathcal{M}}^{{\mathcal{D}}}f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\;\omega)}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}$
(3.16) $\lesssim[\omega]_{{\bf B}_{\frac{1}{3-2t}}}^{\frac{3-2t}{2t-2}}\cdot\left\|f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\;\omega)}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2},$
where in the last estimate, we used the assumption $\omega\in{\bf B}_{\frac{1}{3-2t}}$. Finally for the double summation in (3.2), using the fact that $\{\widetilde{Q}_{4^{tk},i}^{\textnormal{up}}\}_{k\in\mathbb{Z},\;i\geq 1}$ are mutually disjoint, we have
$\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}$ $\lesssim\int_{{\mathbb{D}}}\frac{1}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}dA(z)$
@@
On the other hand side, by the argument in (3.12), we see that for $1\leq k\leq N$,
@@
$\left\|{\mathcal{M}}_{t}^{{\mathcal{D}}}f_{N}\right\|_{L^{1}({\mathbb{D}})}$ $\geq\sum_{k=1}^{N}\int_{D_{k}}{\mathcal{M}}_{t}^{{\mathcal{D}}}f_{N}(z)dA(z)$
$\gtrsim\sum_{k=1}^{N}2^{-k}\cdot 2^{2k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}$
(3.18) $=\sum_{k=1}^{N}2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}.$
@@ 3.3. A two-weight estimate for $\mathcal{M}_{t}^{\mathcal{D}}$
Therefore, using (3.3) with $E_{\alpha}$ replaced by $E_{4^{tk}}\backslash E_{4^{t(k+1)}}$, and $\left\{Q_{4^{tk},i}\right\}_{i\geq 1}$ by $\left\{\widetilde{Q}_{4^{tk},i}\right\}_{i\geq 1}$, respectively, we derive that
@@
$\lesssim\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\frac{\mu(\widetilde{Q}_{4^{tk},i})}{|\widetilde{Q}_{4^{tk},i}|^{tq}}\left(\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}{\mathcal{M}}^{{\mathcal{D}},\sigma}\left(|f|\sigma^{-1}\right)(z)\cdot\sigma^{\frac{1}{p}}(z)\sigma^{\frac{1}{p^{\prime}}}(z)dA(z)\right)^{q}$
$\lesssim\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\frac{\mu(\widetilde{Q}_{4^{tk},i})}{|\widetilde{Q}_{4^{tk},i}|^{tq}}\left(\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)\right)^{\frac{q}{p}}\cdot\left(\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}\sigma(z)dA(z)\right)^{\frac{q}{p^{\prime}}}$
$=\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\frac{\mu(\widetilde{Q}_{4^{tk},i})\sigma^{\frac{q}{p^{\prime}}}(\widetilde{Q}_{4^{tk},i}^{\textrm{up}})}{|\widetilde{Q}_{4^{tk},i}|^{tq}}\cdot\left(\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)\right)^{\frac{q}{p}}$
(3.22) $\leq\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)\right)^{\frac{q}{p}}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\left(\frac{\mu(\widetilde{Q}_{4^{tk},i})\sigma^{\frac{q}{p^{\prime}}}(\widetilde{Q}_{4^{tk},i}^{\textrm{up}})}{|\widetilde{Q}_{4^{tk},i}|^{tq}}\right)^{\frac{p}{p-q}}\right)^{\frac{p-q}{p}}.$
@@
$\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)$ $\lesssim\int_{{\mathbb{D}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)$
@@
$\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\left(\frac{\mu(\widetilde{Q}_{4^{tk},i})\sigma^{\frac{q}{p^{\prime}}}(\widetilde{Q}_{4^{tk},i}^{\textrm{up}})}{|\widetilde{Q}_{4^{tk},i}|^{tq}}\right)^{\frac{p}{p-q}}$ $\simeq\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|\widetilde{Q}_{4^{tk},i}^{\textrm{up}}|\left(\frac{\mu(\widetilde{Q}_{4^{tk},i})\sigma^{\frac{q}{p^{\prime}}}(\widetilde{Q}_{4^{tk},i}^{\textrm{up}})}{|\widetilde{Q}_{4^{tk},i}|^{tq}|\widetilde{Q}_{4^{tk},i}|^{\frac{p-q}{p}}}\right)^{\frac{p}{p-q}}$
(3.24) $\lesssim\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}\phi^{\frac{p}{p-q}}(z)dA(z)\leq\int_{{\mathbb{D}}}\phi^{\frac{p}{p-q}}(z)dA(z).$
@@
where in the last estimate above, we have used the assumption that $\sigma$ is dyadic doubling. Therefore,

### v5

@@ 2. Notations
The Lorentz space $L^{p,r}({\mathbb{D}})$ is defined to be the collection of all measurable functions $f$ such that
$\left\|f\right\|_{L^{p,r}({\mathbb{D}})}:=\left(p\int_{0}^{\infty}t^{r}\left|\left\{z\in{\mathbb{D}}:|f(z)|>\lambda\right\}\right|^{\frac{r}{p}}\frac{d\lambda}{\lambda}\right)^{\frac{1}{r}},\qquad 0<r<+\infty,$
@@ 3.1. Off-critical line estimate
For any $0<\varepsilon\leq 3-2t$, ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{\infty}({\mathbb{D}})\to L^{\frac{1}{2t-2+\varepsilon}}({\mathbb{D}})$ is bounded.
@@
Observe that it can happen that the intersection of the two sets of Carleson boxes $\{Q_{4^{t\ell},i}\}_{i\geq 1}$ and $\{Q_{4^{t(\ell+1)},i}\}_{i\geq 1}$ is not empty. Hence, for each $\ell\in\mathbb{Z}$, define the disjoint union of Carleson boxes
@@ 1. $\bullet$
    $E_{4^{t\ell}}\backslash E_{4^{t(\ell+1)}}\subseteq\bigcup_{i\geq 1}\widetilde{Q}_{4^{t\ell},i}$.
@@ 2. $\bullet$
    $\{\widetilde{Q}^{\textrm{up}}_{4^{t\ell},i}\}_{\ell\in\mathbb{Z},\;i\geq 1}$ are mutually disjoint.
Using now (3.1) with $E_{\alpha}$ replaced by $E_{4^{t\ell}}\backslash E_{4^{t(\ell+1)}}$, and $\{Q_{\alpha,i}\}_{i\geq 1}$ replaced by $\{\widetilde{Q}_{4^{t\ell},i}\}_{i\geq 1}$, respectively, we have
RHS of (3.3) $\leq\sum_{\ell\in\mathbb{Z}}\left(\sum_{i=1}^{\infty}\int_{\widetilde{Q}^{\textnormal{up}}_{4^{t\ell},i}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)\right)^{3-2t-\varepsilon}\cdot\left(\sum_{i=1}^{\infty}|\widetilde{Q}_{4^{t\ell},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}\right)^{2t+\varepsilon-2}$
(3.5) $\leq\left(\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}^{\textnormal{up}}_{4^{t\ell},i}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)\right)^{3-2t-\varepsilon}\cdot\left(\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}|\widetilde{Q}_{4^{t\ell},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}\right)^{2t+\varepsilon-2}.$
Since $\{\widetilde{Q}^{\textnormal{up}}_{4^{t\ell},i}\}_{\ell\in\mathbb{Z},\;i\geq 1}$ are mutually disjoint, therefore, the first double sum in (3.1) is bounded above by
@@
$\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}|\widetilde{Q}_{4^{t\ell},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}$ $\simeq\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}|\widetilde{Q}^{\textnormal{up}}_{4^{t\ell},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}=\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\left|\widetilde{Q}^{\textnormal{up}}_{4^{t\ell},i}\right|\cdot\frac{1}{\left|\widetilde{Q}^{\textnormal{up}}_{4^{t\ell},i}\right|^{\frac{t-1}{2t+\varepsilon-2}}}$
$\simeq\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}^{\textnormal{up}}_{4^{t\ell},i}}\frac{1}{(1-|z|^{2})^{\frac{2(t-1)}{2t+\varepsilon-2}}}dA(z)\lesssim\int_{{\mathbb{D}}}\frac{1}{(1-|z|^{2})^{\frac{2(t-1)}{2t+\varepsilon-2}}}dA(z)$
$\simeq\int_{0}^{1}\frac{1}{(1-r)^{\frac{2(t-1)}{2t+\varepsilon-2}}}dr<+\infty.$
@@ 3.2. Critical line estimate
We first deal with the endpoint $\left(1/p,1/q\right)=(0,2t-2)$.
@@
Sufficiency. Let $\alpha>0$ and denote $E:=\left\{z\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)>\alpha\right\}$. As usual, we decompose $E$ into a union of maximal and mutually disjoint Carleson boxes $\{Q_{i}\}_{i\geq 1}$ satisfying $Q_{i}=Q_{I_{i}}$ for $I_{i}\in{\mathcal{D}}$ and
@@
then for every $0<r<\infty$, the operator ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ extends a bounded operator from $L^{p,r}({\mathbb{D}})$ to $L^{q,r}({\mathbb{D}})$. In particular, ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{p}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$ is bounded.
@@
The first assertion follows from the off-diagonal Marcinkiewicz interpolation theorem [17, Theorem 1.4.19]. The second follows from by letting $r=\infty$ and the fact that $L^{p}({\mathbb{D}})\subset L^{p,\infty}({\mathbb{D}})$. ∎
@@ 2. $\bullet$
Sufficiency. Again, for any $\alpha>0$, we denote $E_{\alpha}:=\left\{z\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)>\alpha\right\}$. Then as usual, one can decompose $E_{\alpha}$ into disjoint union of maximal Carleson boxes $\{Q_{\alpha,i}\}_{i\geq 1}$ enjoying (3.1). Then following the argument in (3.1) with applying Hölder for the conjugate pair $(l,l^{\prime})=\left(\frac{1}{3-2t},\frac{1}{2t-2}\right)$, we have
@@
Then using (3.2) with $E_{\alpha}$ replaced by $E_{4^{t\ell}}\backslash E_{4^{t(\ell+1)}}$, and $\{Q_{\alpha,i}\}_{i\geq 1}$ replaced by $\{\widetilde{Q}_{4^{t\ell},i}\}_{i\geq 1}$, respectively, we have
@@
$\leq\sum_{\ell\in\mathbb{Z}}\left(\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{t\ell}i}^{\textnormal{up}}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t}}\omega(z)dA(z)\right)^{3-2t}\left(\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{t\ell},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}$
$\lesssim\left\|{\mathcal{M}}^{{\mathcal{D}}}f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\;\omega)}\cdot\left(\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{t\ell},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}$
(3.16) $\lesssim[\omega]_{{\bf B}_{\frac{1}{3-2t}}}^{\frac{3-2t}{2t-2}}\cdot\left\|f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\;\omega)}\cdot\left(\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{t\ell},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2},$
where in the last estimate, we used the assumption $\omega\in{\bf B}_{\frac{1}{3-2t}}$. Finally for the double summation in (3.2), using the fact that $\{\widetilde{Q}_{4^{t\ell},i}^{\textnormal{up}}\}_{\ell\in\mathbb{Z},\;i\geq 1}$ are mutually disjoint, we have
$\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{t\ell},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}$ $\lesssim\int_{{\mathbb{D}}}\frac{1}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}dA(z)$
@@
On the other hand side, by the argument in (3.12), we see that for $0\leq k\leq N$,
@@
$\left\|{\mathcal{M}}_{t}^{{\mathcal{D}}}f_{N}\right\|_{L^{1}({\mathbb{D}})}$ $\geq\sum_{k=0}^{N}\int_{D_{k}}{\mathcal{M}}_{t}^{{\mathcal{D}}}f_{N}(z)dA(z)$
$\gtrsim\sum_{k=0}^{N}2^{-k}\cdot 2^{2k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}$
(3.18) $=\sum_{k=0}^{N}2^{k}\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}.$
@@ 3.3. A two-weight estimate for $\mathcal{M}_{t}^{\mathcal{D}}$
Therefore, using (3.3) with $E_{\alpha}$ replaced by $E_{4^{t\ell}}\backslash E_{4^{t(\ell+1)}}$, and $\left\{Q_{\alpha,i}\right\}_{i\geq 1}$ by $\left\{\widetilde{Q}_{4^{t\ell},i}\right\}_{i\geq 1}$, respectively, we derive that
@@
$\lesssim\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\frac{\mu(\widetilde{Q}_{4^{t\ell},i})}{|\widetilde{Q}_{4^{t\ell},i}|^{tq}}\left(\int_{\widetilde{Q}_{4^{t\ell},i}^{\textrm{up}}}{\mathcal{M}}^{{\mathcal{D}},\sigma}\left(|f|\sigma^{-1}\right)(z)\cdot\sigma^{\frac{1}{p}}(z)\sigma^{\frac{1}{p^{\prime}}}(z)dA(z)\right)^{q}$
$\lesssim\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\frac{\mu(\widetilde{Q}_{4^{t\ell},i})}{|\widetilde{Q}_{4^{t\ell},i}|^{tq}}\left(\int_{\widetilde{Q}_{4^{t\ell},i}^{\textrm{up}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)\right)^{\frac{q}{p}}\cdot\left(\int_{\widetilde{Q}_{4^{t\ell},i}^{\textrm{up}}}\sigma(z)dA(z)\right)^{\frac{q}{p^{\prime}}}$
$=\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\frac{\mu(\widetilde{Q}_{4^{t\ell},i})\sigma^{\frac{q}{p^{\prime}}}(\widetilde{Q}_{4^{t\ell},i}^{\textrm{up}})}{|\widetilde{Q}_{4^{t\ell},i}|^{tq}}\cdot\left(\int_{\widetilde{Q}_{4^{t\ell},i}^{\textrm{up}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)\right)^{\frac{q}{p}}$
(3.22) $\leq\left(\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{t\ell},i}^{\textrm{up}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)\right)^{\frac{q}{p}}\cdot\left(\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\left(\frac{\mu(\widetilde{Q}_{4^{t\ell},i})\sigma^{\frac{q}{p^{\prime}}}(\widetilde{Q}_{4^{t\ell},i}^{\textrm{up}})}{|\widetilde{Q}_{4^{t\ell},i}|^{tq}}\right)^{\frac{p}{p-q}}\right)^{\frac{p-q}{p}}.$
@@
$\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{t\ell},i}^{\textrm{up}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)$ $\lesssim\int_{{\mathbb{D}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)$
@@
$\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\left(\frac{\mu(\widetilde{Q}_{4^{t\ell},i})\sigma^{\frac{q}{p^{\prime}}}(\widetilde{Q}_{4^{t\ell},i}^{\textrm{up}})}{|\widetilde{Q}_{4^{t\ell},i}|^{tq}}\right)^{\frac{p}{p-q}}$ $\simeq\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}|\widetilde{Q}_{4^{t\ell},i}^{\textrm{up}}|\left(\frac{\mu(\widetilde{Q}_{4^{t\ell},i})\sigma^{\frac{q}{p^{\prime}}}(\widetilde{Q}_{4^{t\ell},i}^{\textrm{up}})}{|\widetilde{Q}_{4^{t\ell},i}|^{tq}|\widetilde{Q}_{4^{t\ell},i}|^{\frac{p-q}{p}}}\right)^{\frac{p}{p-q}}$
(3.24) $\lesssim\sum_{\ell\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{t\ell},i}^{\textrm{up}}}\phi^{\frac{p}{p-q}}(z)dA(z)\leq\int_{{\mathbb{D}}}\phi^{\frac{p}{p-q}}(z)dA(z).$
@@
where in the last estimate above, we have used the assumption that $\sigma\in{\bf B}_{\infty}({\mathcal{D}})$. Therefore,